OpenStudy (detectivefluffy):

Find the least common multiple of these two expressions: 4v^3y^5 and 18u^4v^2y^6

OpenStudy (jdoe0001):

hint: \(\bf \begin{cases} 4v^3y^5\implies &2\cdot {\color{blue}{ 2}}\cdot {\color{blue}{ v\cdot v}}\cdot v\cdot {\color{blue}{ y\cdot y\cdot y\cdot y\cdot y }} \\ \quad \\ \quad \\ 18u^4v^2y^6\implies &3\cdot 3\cdot {\color{blue}{ 2}}\cdot u\cdot u\cdot u\cdot u\cdot {\color{blue}{ v\cdot v}} \cdot {\color{blue}{ y\cdot y\cdot y\cdot y\cdot y }}\cdot y \end{cases}\)

OpenStudy (detectivefluffy):

So would the answer be 36v^3u^4y^6?

OpenStudy (jdoe0001):

well, you're being asked for the LCM, or LCD or GCF, usually goes by any of those names so.... what factors are COMMON to both terms?

OpenStudy (detectivefluffy):

2v^2y^5